OpenStudy (anonymous):

for the quadratic function f(x) = (x - (1/3))62 + 2 in standard form, identify the domain, range, vertex, and axis of symmetry of its graph. i got domain is all real numbers. vertex: ((1/3), 2); axis of symmetry is x = (1/3)

OpenStudy (anonymous):

range: 

OpenStudy (jdoe0001):

can you see where the vertex is?

OpenStudy (anonymous):

the quadratic function is suppose to be f(x) = (x-(1/3))^2 + 2

OpenStudy (anonymous):

vertex is (1/3), 2?

OpenStudy (jdoe0001):

yes, and the parabola is going upwards, leading coefficient is positive, so it has a minimum, thus the range is correct \(\bf \boldsymbol{ [2, +\infty)}\)

OpenStudy (anonymous):

Thank you!

OpenStudy (jdoe0001):

the domain will be \(\bf x \in \mathbb{R}\)

OpenStudy (jdoe0001):